# Optimizing for ROC Curves on Class-Imbalanced Data by Training over a Family of Loss Functions

### v1

@@ 5 Optimizing for ROC Curves via Loss Conditional Training
We saw in Section 4.3 that training with different combinations of hyperparameters corresponds to optimizing for different tradeoffs of TPR and FPR. Therefore, training one model over a range of hyperparameter values is a proxy for optimizing over a range of the ROC curve. Using this intuition, we design a system that optimizes one model over a range of VS loss hyperparameter values via loss conditional training (LCT).
@@ 5.1 Loss Conditional Training (LCT)
Let $\boldsymbol{\lambda}$ be a vector which parameterizes a loss function. For example, we can parameterize VS loss by $\boldsymbol{\lambda}=(\Omega,\gamma,\tau)$. Then let $\mathcal{L}(\cdot,\cdot,\boldsymbol{\lambda})$ be the family of loss functions parameterized by $\boldsymbol{\lambda}\in\boldsymbol{\Lambda}\subseteq\mathbb{R}^{d_{\boldsymbol{\lambda}}}$ where $d_{\boldsymbol{\lambda}}$ is the length of $\boldsymbol{\lambda}$. In our example, $\mathcal{L}(\cdot,\cdot,\boldsymbol{\lambda})$ is the set of all possible VS loss functions obtained by different parameter values for $\Omega,\gamma,\tau$.
Normally, training finds the weights $\boldsymbol{\theta}$ on a model $f$ which minimize a single loss function $\mathcal{L}(\cdot,\cdot,\boldsymbol{\lambda})$ (i.e., we optimize for a single combination of $\boldsymbol{\lambda}_{0}=(\Omega_{0},\gamma_{0},\tau_{0})$ values in VS loss) as shown below.
$\boldsymbol{\theta}^{*}_{\lambda_{0}}$ $=\operatorname*{arg\,min}_{\boldsymbol{\theta}}\mathbb{E}_{(\mathbf{x},y)\sim D}\mathcal{L}(y,f(\mathbf{x},\boldsymbol{\theta}),\boldsymbol{\lambda}_{0})$ (7)
LCT instead optimizes over a distribution of $\boldsymbol{\lambda}$ values $P_{\boldsymbol{\Lambda}}$ as shown below.
$\boldsymbol{\theta}^{*}$ $=\operatorname*{arg\,min}_{\boldsymbol{\theta}}\mathbb{E}_{\boldsymbol{\lambda}\sim P_{\boldsymbol{\Lambda}}}\mathbb{E}_{(\mathbf{x},y)\sim D}\mathcal{L}(y,f(\mathbf{x},\boldsymbol{\theta},\boldsymbol{\lambda}),\boldsymbol{\lambda})$ (8)
LCT is implemented on Deep Neural Network (DNN) predictors by augmenting the network to take an additional input vector $\boldsymbol{\lambda}$ along with each data sample $\mathbf{x}$. During training, $\boldsymbol{\lambda}$ is sampled from $\boldsymbol{\Lambda}$ with each sample and is used in two ways: 1) as an additional input to the network and 2) in the loss function for the sample. During inference, the model takes as input a $\boldsymbol{\lambda}$ along with the sample and outputs the appropriate prediction.
In order to condition the model on $\boldsymbol{\lambda}$, the DNN is augmented with Feature-wise Linear Modulation (FiLM) layers (Perez et al., 2018). These are small neural networks that take the conditioning parameter $\boldsymbol{\lambda}$ as input and output a $\boldsymbol{\mu}$ and $\boldsymbol{\sigma}$, used to modulate the activations channel-wise based on the value of $\boldsymbol{\lambda}$. Specifically, suppose a layer in the DNN has an activation map of size $H\times W\times C$. In LCT, we transform each activation $\mathbf{f}\in\mathbb{R}^{C}$ by $\boldsymbol{\mu}$ and $\boldsymbol{\sigma}$ as follows: $\mathbf{\tilde{f}}=\boldsymbol{\sigma}*\mathbf{f}+\boldsymbol{\mu}$ where both $\boldsymbol{\mu}$ and $\boldsymbol{\sigma}$ are vectors of size $C$, and “***” stands for element-wise multiplication.
@@ 5.2 Details about implementing LCT for classification
We define our family of loss functions using VS loss. We try several choices for $\boldsymbol{\lambda}$, including $\boldsymbol{\lambda}=\Omega,\boldsymbol{\lambda}=\gamma,\boldsymbol{\lambda}=\tau$ and $\boldsymbol{\lambda}=(\Omega,\gamma,\tau)$. In the first three cases, we set the hyperparameters excluded from $\boldsymbol{\lambda}$ to constants (e.g., when $\boldsymbol{\lambda}=\tau$, $\Omega$ and $\gamma$ are set to constants). We find that $\boldsymbol{\lambda}=\tau$ is especially strong at improving the average performance and reducing the variance in performance (Figure 8), which is supported by our insights in Figure 3 and Section 4.3. Thus, for most of our experiments, we use $\boldsymbol{\lambda}=\tau$.
For each mini-batch, we draw one $\boldsymbol{\lambda}$. This is done by independently sampling each hyperparameter in $\boldsymbol{\lambda}$ from a distribution that has a linear density over range $[a,b]$. In this distribution, the user specifies $a,b$, and the height of the probability density function (pdf) at $b$, $h_{b}$. The height at $a$, $h_{a}$, is then found to ensure the area under the pdf is 1. Unlike the triangular distribution, this general linear distribution does not require the pdf to be 0 at either endpoint and, unlike the uniform function, it does not require the slope of the pdf to be 0.555Appendix E contains more details about the linear distribution. We use this general distribution as a way to experiment with different distributions.
Although the value of $\boldsymbol{\lambda}$ at inference time affects the TPR and FPR when $t=0.5$, we observe that the ROC curves are almost identical for all values of $\boldsymbol{\lambda}$ in the range the model was trained on (see Appendix D). Thus, we evaluate LCT models at one $\boldsymbol{\lambda}$ and find their ROC curve from this output.
@@ 6.1 Experimental setup
Methods. For each dataset and imbalance ratio $\beta$, we train 48 models with the regular VS loss, varying their hyperparameter values. We also train 48 models with LCT applied to VS loss where $\boldsymbol{\lambda}=\tau$ (VS loss + LCT). We use $\Omega=0.5,0.7,0.9,0.99$, $\gamma=0.0,0.2,0.4$ for both methods (note that we set $\Omega,\gamma$ to constants when $\boldsymbol{\lambda}=\tau$ in LCT). We use $\tau=0,1,2,3$ for VS loss, and $[a,b]=[0,3]$ with $h_{b}=0.0,0.15,0.33,0.66$ for VS loss + LCT. We evaluate with $\boldsymbol{\lambda}=\tau=3$. To apply LCT, we augment the networks with one FiLM block after the final convolutional layer and before the linear layer. This block is comprised of two linear layers with 128 hidden units. Specifically, the first layer takes one input $\boldsymbol{\lambda}$ and outputs 128 hidden values and the second layer outputs 64 values which are used to modulate the 64 channels of convolutional activations (total of $1*128+128*64=8320$ parameters if $\boldsymbol{\lambda}$ is a scalar).
@@
Datasets. We experiment on both toy datasets derived from CIFAR10/100 and more realistic datasets derived from Kaggle competitions. CIFAR10 cat/dog consists of the cat and dog classes from the CIFAR10 dataset. We also show results for all pairs of CIFAR10 classes in Section 6.3, but find the cat/dog pair is particularly well-suited for additional experiments since it is a challenging classification problem. CIFAR100 household electronics vs. furniture was proposed as a binary dataset with imbalance by (Wang et al., 2016). Each class contains 5 classes from CIFAR100: electronics contains clock, computer keyboard, lamp, telephone, and television while furniture contains bed, chair, couch, table, and wardrobe. For all CIFAR experiments, we split the data according to their train and test splits. Kaggle Dogs vs. Cats contains 25,000 images of dogs and cats from the Petfinder website (Cukierski, 2013). We split the data so that each class has 1,000 validation samples (a 92/8 train/validation split). Finally, SIIM-ISIC Melanoma is a classification dataset, which was designed by the Society for Imaging Informatics in Medicine (SIIM) and the International Skin Imaging Collaboration (ISIC) for a Kaggle competition (Zawacki et al., 2020). This is a binary dataset where 8.8% of the samples belong to the positive (melanoma) class (i.e., $\beta=12$). We follow the procedure of (Fang et al., 2023) and combine the 33,126 and 25,331 images from 2020 and 2019 respectively and split the data into an 80/20 train/validation split. Table 9 outlines the sizes of these. We subsample the minority class to obtain various imbalance ratios $\beta$. Specifically, we test the CIFAR datasets at $\beta=10,50,100,200$, the Kaggle datasets at $\beta=100,200$.
@@
Model architectures and training procedure. We use two model architectures and training procedures, following the literature for each type of dataset. For CIFAR10/100 data, we follow (Kini et al., 2021) and use a ResNet-32 model architecture. We find that LCT takes longer to train, so we train for 500 epochs instead of 200 (note: training the baseline models longer does not significantly alter performance and we do this for consistency). We train these models from scratch using an initial learning rate of 0.1 and decrease this to 0.01 and 0.001 at 400 and 450 epochs respectively. For the Kaggle datasets, we follow (Shwartz-Ziv et al., 2023) and use ResNext50-32x4d with weights pre-trained on ImageNet. We train the Melanoma dataset for 10 epochs, following the learning rates and weight decays of Fang et al. (2023) and train the Kaggle Dogs vs. cats models for 30 epochs, following the learning rates and weight decays of other finetuning tasks in (Fang et al., 2023). For all experiments, we use a batch size of 128 and gradient clipping with a maximum norm equal to 0.5. Additionally, except for Figure 7, we use Stochastic Gradient Descent (SGD) optimization with momentum=0.9. We run all experiments on A5000 GPUs.
@@
Figure 5: Distribution of ROC curves of models trained with (red) and without (blue) LCT on CIFAR datasets at four different imbalance ratios $\beta$. Solid, dashed, and dotted-dashed curves are the mean, minimum, and maximum of the ROC curves respectively. The shaded region is one standard deviation away from the mean. Datasets are tested with $\beta=10,50,100,200$. At high imbalance ratios, LCT consistently improves the mean, max, and min of the ROC curves.
@@
Figure 6: Distribution of ROC curves of models trained with (red) and without (blue) LCT on Kaggle Dogs vs. Cats and SIIM-ISIC Melanoma datasets at two different imbalance ratios $\beta$. Solid, dashed, and dotted-dashed curves are the mean, minimum, and maximum of the ROC curves respectively. The shaded region is one standard deviation away from the mean. Results are shown on a log scale because all TPR values are compressed towards 1. LCT significantly improves performance on Melanoma, especially at high TPRs.
@@ 6.2 VS loss with and without LCT
Figure 5 shows that at moderate imbalance ratios ($\beta=10$), there is very little variation in the ROC curves obtained by models trained with different hyperparameters. Here the two methods are essentially tied in their performance. However, at high imbalance ratios $\beta=100,200$ (Figures 5 and 6), the variance in ROC curves becomes much more drastic. Here LCT consistently outperforms the Baseline method. Specifically, LCT improves the max and mean of the ROC curves in all cases, with a large gap in improvement for Melanoma at $\beta=100,200$. Additionally, LCT significantly decreases the variation in the results. Notice how the highlighted bands, representing the standard deviation, are much narrower for LCT (red) than Baseline (blue). Furthermore, the LCT’s minimum (red dashed) curves are much higher Baseline’s (blue dashed). We also see this in the distribution of AUC values shown in Figure 1. This reduction in variance indicates that LCT is more robust to hyperparameter changes and hence requires less tuning.
@@ 6.3 Evaluation over all CIFAR10 pairs
For a more comprehensive evaluation, we compare the results of VS loss with and without LCT on all 45 pairs of classes chosen from the 10 CIFAR10 classes. We train each method on each of these binary datasets using 48 different hyperparameter values. We then analyze the aggregated performance of these models by AUC. For each binary dataset, we find the mean, maximum, minimum, and standard deviation of the AUCs over the hyperparameter values. We then compare Baseline (VS loss) and LCT (VS loss + LCT) for each of these metrics. For example, let $a^{max}_{k,\text{Baseline}}$ and $a^{max}_{k,\text{LCT}}$ be the maximum AUCs achieved on dataset $k$ by LCT and Baseline respectively. Then we obtain 45 of these values for each method (one for each dataset $k$). We then analyze these using a paired t-test and report results in Table 1. The table shows that LCT improves performance, increasing the min and mean AUC in all cases, as well as the max AUC in most cases. It also drastically reduces the standard deviation of AUC values. This shows that LCT significantly outperforms Baseline on a comprehensive set of datasets.
@@
Previous work has shown that Sharpness Aware Minimization (SAM) is effective at optimizing imbalanced datasets in the multi-class setting (Foret et al., 2021; Rangwani et al., 2022). SAM searches for parameters that lie in neighborhoods of uniformly low loss and uses a hyperparameter $\rho$ to define the size of these neighborhoods. Rangwani et al. (2022) shows that SAM can effectively escape saddle points in imbalanced settings by using large values for $\rho$ (i.e., $\rho=0.5,0.8$). We, however, find that using SAM in this way does not translate to improved performance on a binary dataset. In Figure 7, we compare the performance of the models across several hyperparameter values $\rho$. We find that, unlike the multi-class case, in the binary case the best $\rho$ is quite small (i.e., $\rho=0.1$). While training with SAM and $\rho=0.1$ can produce some very strong results, this also exacerbates the amount of variance in the performance. LCT with SGD gives models that are better on average and require less tuning. Note that larger $\rho$ values are detrimental to training in this setting and lead to models that are effectively naive classifiers. We also tried combining LCT and SAM with $\rho=0.1$ and found this to be less effective than training with LCT and SGD.
Figure 7: Effects of using SAM optimization. Each boxplot contains results for 48 models trained with different hyperparameter values. Models were trained on CIFAR10 cat/dog $\beta=100$. In the binary case, SAM fails with large $\rho$. Although some models trained with SAM and small $\rho$ achieve strong performance, SAM exacerbates the variance of the results.

### v2

@@ 5 Optimizing for ROC Curves via Loss Conditional Training
Recall that ROC curves show the performance of the model in terms of TPR for all possible FPRs and vice versa. We saw in Section 4.3 that training with different combinations of hyperparameters corresponds to optimizing for different tradeoffs of TPR and FPR. Therefore, training one model over a range of hyperparameter values is a proxy for optimizing over a range of the ROC curve. Using this intuition, we design a system which optimizes one model over a range of VS loss hyperparameter values via loss conditional training (LCT).
@@ 5.1 Loss Conditional Training (LCT)
Let $\bm{\lambda}$ be a vector which parameterizes a loss function. For example, we can parameterize VS loss by $\bm{\lambda}=(\Omega,\gamma,\tau)$. Then let $\mathcal{L}(\cdot,\cdot,\bm{\lambda})$ be the family of loss functions parameterized by $\bm{\lambda}\in\bm{\Lambda}\subseteq\mathbb{R}^{d_{\bm{\lambda}}}$ where $d_{\bm{\lambda}}$ is the length of $\bm{\lambda}$. In our example, $\mathcal{L}(\cdot,\cdot,\bm{\lambda})$ is the set of all possible VS loss functions obtained by different parameter values for $\Omega,\gamma,\tau$.
Normally, training finds the weights $\bm{\theta}$ on a model $f$ which minimize a single loss function $\mathcal{L}(\cdot,\cdot,\bm{\lambda})$ (i.e., we optimize for a single combination of $\bm{\lambda}_{0}=(\Omega_{0},\gamma_{0},\tau_{0})$ values in VS loss) as shown below.
$\bm{\theta}^{*}_{\lambda_{0}}$ $=\operatorname*{arg\,min}_{\bm{\theta}}\mathbb{E}_{(\mathbf{x},y)\sim D}\mathcal{L}(y,f(\mathbf{x},\bm{\theta}),\bm{\lambda}_{0})$ (7)
LCT instead optimizes over a distribution of $\bm{\lambda}$ values $P_{\bm{\Lambda}}$ as shown below.
$\bm{\theta}^{*}$ $=\operatorname*{arg\,min}_{\bm{\theta}}\mathbb{E}_{\bm{\lambda}\sim P_{\bm{\Lambda}}}\mathbb{E}_{(\mathbf{x},y)\sim D}\mathcal{L}(y,f(\mathbf{x},\bm{\theta},\bm{\lambda}),\bm{\lambda})$ (8)
LCT is implemented on Deep Neural Network (DNN) predictors by augmenting the network to take an additional input vector $\bm{\lambda}$ along with each data sample $\mathbf{x}$. During training, $\bm{\lambda}$ is sampled from $\bm{\Lambda}$ with each sample and is used in two ways: 1) as an additional input to the network and 2) in the loss function for the sample. During inference, the model takes as input a $\bm{\lambda}$ along with the sample and outputs the appropriate prediction.
In order to condition the model on $\bm{\lambda}$, the DNN is augmented with Feature-wise Linear Modulation (FiLM) layers Perez et al. (2018). These are small neural networks that take the conditioning parameter $\bm{\lambda}$ as input and output a $\bm{\mu}$ and $\bm{\sigma}$, used to modulate the activations channel-wise based on the value of $\bm{\lambda}$. Specifically, suppose a layer in the DNN has an activation map of size $H\times W\times C$. In LCT, we transform each activation $\mathbf{f}\in\mathbb{R}^{C}$ by $\bm{\mu}$ and $\bm{\sigma}$ as follows: $\mathbf{\tilde{f}}=\bm{\sigma}*\mathbf{f}+\bm{\mu}$ where both $\bm{\mu}$ and $\bm{\sigma}$ are vectors of size $C$, and “*” stands for element-wise multiplication.
@@ 5.2 Details about implementing LCT for classification
We define our family of loss functions $\bm{\lambda}$ using VS loss. We try several choices for $\bm{\lambda}$, including $\bm{\lambda}=\Omega,\bm{\lambda}=\gamma,\bm{\lambda}=\tau$ and $\bm{\lambda}=(\Omega,\gamma,\tau)$. In the first three cases, we set the hyperparameters excluded from $\bm{\lambda}$ to constants (as is done in regular VS loss). We find that $\bm{\lambda}=\tau$ is especially strong at improving the average performance and reducing the variance in performance (Figure 8), which is supported by our insights in Figure 3 and Section 4.3. Thus, for most of our experiments, we use $\bm{\lambda}=\tau$.
For each mini-batch, we draw one $\bm{\lambda}$. This is done by independently sampling each hyperparameter in $\bm{\lambda}$ from a distribution that has a linear density over range $[a,b]$. In this distribution, the user specifies $a,b$, and the height of the probability density function (pdf) at $b$, $h_{b}$. The height at $a$, $h_{a}$, is then found to ensure the area under the pdf is 1. Unlike the triangular distribution, this general linear distribution does not require the pdf to be 0 at either endpoint and, unlike the uniform function, it does not require the slope of the pdf to be 0.444Appendix D contains more details about the linear distribution. We use this general distribution as a way to experiment with different distributions.
Although the value of $\bm{\lambda}$ at inference time affects the TPR and FPR when $t=0.5$, we observe that the ROC curves are almost identical for all values of $\bm{\lambda}$ in the range the model was trained on (see Appendix C). Thus, we evaluate LCT models at one $\bm{\lambda}$ and find their ROC curve from this output.
@@ 6.1 Experimental setup
Methods. For each dataset and imbalance ratio $\beta$, we train 48 models with the regular VS loss, varying their hyperparameter values. We also train 48 models with LCT applied to VS loss where $\bm{\lambda}=\tau$ (VS Loss + LCT). We use $\Omega=0.5,0.7,0.9,0.99$, $\gamma=0.0,0.2,0.4$ for both methods (note that we set $\Omega,\gamma$ to constants when $\bm{\lambda}=\tau$ in LCT). We use $\tau=0,1,2,3$ for VS loss, and $[a,b]=[0,3]$ with $h_{b}=0.0,0.15,0.33,0.66$ for VS loss + LCT. We evaluate with $\bm{\lambda}=\tau=3$. To apply LCT, we augment the networks with one FiLM block after the final convolutional layer and before the linear layer. This block is comprised of two linear layers with 128 hidden units. Specifically, the first layer takes one input $\bm{\lambda}$ and outputs 128 hidden values and the second layer outputs 64 values which are used to modulate the 64 channels of convolutional activations (total of $1*128+128*64=8320$ parameters if $\bm{\lambda}$ is a scalar).
@@
Datasets. We experiment on both toy datasets derived from CIFAR10/100 and more realistic datasets derived from Kaggle competitions. CIFAR10 cat/dog consists of the cat and dog classes from the CIFAR10 dataset. We also show results for all pairs of CIFAR10 classes in Section 6.3, but find the cat/dog pair is particularly well-suited for additional experiments since it is a challenging classification problem. CIFAR100 household electronics vs. furniture was proposed as a binary dataset with imbalance by Wang et al. (2016). Each class contains 5 classes from CIFAR100: electronics contains clock, computer keyboard, lamp, telephone and television, and furniture contains bed, chair, couch, table and wardrobe. For all CIFAR experiments, we split the data according to their train and test splits. Kaggle Dogs vs. Cats contains 25,000 images of dogs and cats from the Petfinder website Cukierski (2013). We split the data so that each class has 1,000 validation samples (a 92/8 train/validation split). Finally, SIIM-ISIC Melanoma is a classification dataset, which was designed by the Society for Imaging Informatics in Medicine (SIIM) and the International Skin Imaging Collaboration (ISIC) for a Kaggle competition Zawacki et al. (2020). This is a binary dataset where 8.8% of the samples belong to the positive (melanoma) class (i.e., $\beta=12$). We follow the procedure of Fang et al. (2023) and combine the 33,126 and 25,331 images from 2020 and 2019 respectively and split the data into an 80/20 train/validation split. Table 9 outlines the sizes of these. We subsample the minority class to obtain various imbalance ratios $\beta$. Specifically, we test the CIFAR datasets at $\beta=10,50,100,200$, the Kaggle datasets at $\beta=100,200$.
@@
Model architectures and training procedure. We use two model architectures and training procedures, following the literature for each type of dataset. For CIFAR10/100 data, we follow Kini et al. (2021) and use a ResNet-32 model architecture. We find that LCT takes longer to train, so we train for 500 epochs instead of 200 (note: training the baseline models longer does not significantly alter performance and we do this for consistency). We train these models from scratch using an initial learning rate of 0.1 and decreasing this to $1e^{3}$ and $1e^{5}$ at 400 and 450 epochs respectively. For the Kaggle datasets, we follow Shwartz-Ziv et al. (2023) and use ResNext50-32x4d with weights pre-trained on ImageNet. We train the Melanoma dataset for 10 epochs, following the learning rates and weight decays of Fang et al. (2023) and train the Kaggle Dogs vs. cats models for 30 epochs, following the learning rates and weight decays of other finetuning tasks in Fang et al. (2023). For all experiments we use a batch size of 128 and gradient clipping with maximum norm equal to 0.5. Additionally, with the exception of Figure 8, we use Stochastic Gradient Descent (SGD) optimization with momentum=0.9. We run all experiments on A5000 GPUs.
@@
Figure 5: Distribution of ROC curves of models trained with (red) and without (blue) LCT on CIFAR datasets at four different imbalance ratios $\beta$. Solid, dashed, and dotted-dashed curves are mean, minimum, and maximum of the ROC curves respectively. Shaded region is one standard deviation away from the mean. Datasets are tested with $\beta=10,50,100,200$. At high imbalance ratios, LCT consistently improves the mean, max, and min of the ROC curves.
@@
Figure 6: Distribution of ROC curves of models trained with (red) and without (blue) LCT on Kaggle Dogs vs. Cats and SIIM-ISIC Melanoma datasets at two different imbalance ratios $\beta$. Solid, dashed, and dotted-dashed curves are mean, minimum, and maximum of the ROC curves respectively. Shaded region is one standard deviation away from the mean. Results are shown on log scale because all TPR values are compressed towards 1. LCT significantly improves performance on Melanoma, especially at high TPRs.
@@ 6.2 VS Loss with and without LCT
Figure 5 shows that at moderate imbalance ratios ($\beta=10$), there is very little variation in the ROC curves obtained by models trained with different hyperparameters. Here the two methods are essentially tied in their performance. However, at high imbalance ratios $\beta=100,200$ (Figures 5 and 6), the variance in ROC curves becomes much more drastic. Here LCT consistently outperforms the Baseline method. Specifically, LCT improves or ties the max and mean of the ROC curves in all cases, with a large gap in improvement for Melanoma at $\beta=100,200$.
@@ 6.3 Evaluation over all CIFAR10 pairs
For a more comprehensive evaluation, we compare the results of VS loss with and without LCT on all 45 pairs of classes chosen from the 10 CIFAR10 classes. We train each method on each of these binary datasets using 48 different hyperparameter values. We then analyze the aggregated performance of these models by AUC. For each binary dataset, we find the mean, maximum, minimum, and standard deviation of the AUCs over the hyperparameter values. We then compare Baseline (VS loss) and LCT (VS loss + LCT) for each of these metrics. For example, let $a^{max}_{k,\text{Baseline}}$ and $a^{max}_{k,\text{LCT}}$ be the maximum AUCs achieved on dataset $k$ by LCT and Baseline respectively. Then we obtain 45 of these values for each method (one for each dataset $k$). We then analyze these using a paired t-test and report results in Table 1. The table shows that LCT improves performance, increasing the min and mean AUC in nearly all cases, as well as the max AUC in most cases. It also reduces the standard deviation of AUC values. This shows that LCT significantly outperforms Baseline on a comprehensive set of datasets.
@@
Figure 7: Effects of using SAM optimization. Each boxplot contains results for 48 models trained with different hyperparameter values. Models were trained on CIFAR10 cat/dog $\beta=100$.
Figure 8: LCT with different $\bm{\lambda}$. Each boxplot contains results for 27 models with different hyperparameter values. All models were trained on CIFAR10 cat/dog with $\beta=100$.
@@
Previous work has shown that Sharpness Aware Minimization (SAM) is effective at optimizing imbalanced datasets in the multi-class setting Foret et al. (2021); Rangwani et al. (2022). SAM searches for parameters that lie in neighborhoods of uniformly low loss and usees a hyperparameter $\rho$ to define the size of these neighborhoods. Rangwani et al. (2022) shows that SAM can effectively escape saddle points in imbalanced settings by using large values for $\rho$ (i.e., $\rho=0.5,0.8$). We, however, find that using SAM in this way does not translate to improved performance on a binary dataset. In Figure 8, we compare the performance of the models across several hyperparameter values $\rho$. We find that, unlike the multi-class case, in the binary case the best $\rho$ is quite small (i.e., $\rho=0.1$). While training with SAM and $\rho=0.1$ can produce some very strong results, this also exacerbates the amount of variance in the performance. LCT with SGD gives models which are better on average and require less tuning. Note that larger $\rho$ values are absolutely detrimental to training in this setting and lead to models which are effectively naive classifiers. We also tried combining LCT and SAM with $\rho=0.1$ and found this to be less effective than training with LCT and SGD.